In order to determine whether or not there is a significant difference between the hourly wages of two companies, the following data have been accumulated.

Company 1 Company 2
n1 =58 n2 =64
1 = $10.95 2 = $10.00
S1= $1.81 S2= $1.50

​The point estimate of the difference between the means (Company 1 – Company 2) is _____.
Answer #1

The point estimate of the difference between the means (Company 1 - Company 2) is: X1 - X2= 10.95-10.00 = $0.95 The point est